OpenStudy (anonymous):

What is the derivative of (2x-x^2)^.5 Please simplify

OpenStudy (anonymous):

you need to use the chain rule...

OpenStudy (anonymous):

Yes, \[\frac{ 1 }{ 2 }(2x-x ^{2})^{-.5}(2)(-2x)\]

OpenStudy (anonymous):

@utkFRESH

zepdrix (zepdrix):

Hmm looks very close anthony. Your chain is a little bit off though:\[\Large\rm \frac{d}{dx}(2x-x^2)=(2-2x)\]You have an extra set of brackets that shouldn't be there.

OpenStudy (anonymous):

Thank you very much @zepdrix I see my error now.